# thanks: Corresponding author

### v2

@@ Abstract
We propose the use of the “spin-opstring”, derived from Stochastic Series Expansion Quantum Monte Carlo (QMC) simulations as machine learning input data. It offers a compact, memory-efficient representation of QMC simulation cells, combining the initial state with an operator string that encodes the state’s evolution through imaginary time. Using supervised machine learning, we demonstrate the input’s effectiveness in capturing both conventional and topological phase transitions. Additionally, we conduct a regression task to predict superfluid density, which reflects non-local properties of the quantum system, and achieve good accuracy. We also demonstrate the capability of spin-opstring data in transfer learning by training models on one quantum system and successfully predicting on another, as well as showing that models trained on smaller system sizes generalize well to larger ones. Finally, using two state-of-the-art interpretability techniques, Layer-wise Relevance Propagation and SHapley Additive exPlanations, we show that the machine learning models learn and rely on physically meaningful features from the input data. Together, these findings establish the spin-opstring as a broadly-applicable and interpretable input format for machine learning in quantum many-body physics.
@@ I Introduction
We further explored the application of spin-opstring data for transfer learning tasks. First, we applied our ML model trained on data from the $t_{2}-V_{1}$ model to successfully predict a phase transition in the $t_{1}-V_{1}$ model. Next, we were able to predict phase transitions in the interacting 1D SSH model using a ML model trained in the non-interacting case. Additionally, we showed that it is possible to generalize a ML model trained with spin-opstring data from a smaller system to larger systems. To better understand the benefit of using the full spin-opstring instead of just the spin-state or operator string alone, we discuss some examples, where the full input proves indispensable.
Having established the usefulness of this input datatype, we turn our focus to the interpretation of the decision-making process in the trained ML models. It is important to mention here that the majority of studies dealing with interpretability put their emphasis on spin states as the input data [24, 25, 26]. The suitability of alternative input representations in such analyses remain mostly unexplored [27, 28]. To the best of our knowledge, our study marks the first instance of interpreting ML models trained on simulation cell data. To this end, we utilize two well-known interpretability methods, Layer-wise Relevance Propagation (LRP) [29] and SHapley Additive exPlanations (SHAP) [30] to understand which features in the data are most important to the model’s predictions. We developed a strategy to extract meaningful insights, based on different physical operators, from the large-dimensional feature scores obtained directly from LRP and SHAP. In the classification, regression and transfer learning tasks, we find that ML models trained with spin-opstring data are able to identify key physical features from each phase, showcasing their ability to learn and utilize essential quantum characteristics.
@@ II Proposal
Figure 1: Schematic diagram of the simulation cell of SSE QMC algorithm for a spin-$1/2$ system with six sites. Here $i$ signifies the lattice site number and $\sigma(i)$ represents the initial/final spin state. At each level $p$ of the simulation cell, the spin state is depicted by open (+1)/filled (-1) circles. The diagonal or off-diagonal operator at each level acting on a bond is represented by open or filled bars respectively. The information in the operator string is compactly codified in $S(p)$ ($a(p)$ and $b(p)$ are defined in the text). The “spin-opstring” for the above simulation cell is constructed by combining $\sigma(i)$ and $S(p)$ together.
@@
$|\alpha(p)\rangle=\prod\limits_{i=1}^{p}H_{a(i),b(i)}|\alpha\rangle,|\alpha(0)\rangle=|\alpha\rangle,$ (7)
@@ III Method
For the phase classification in the 2D $t_{2}-V_{1}$ model, we have used a fully-connected ML model with three layers: the first input layer consisting of a number of neurons equal to the spin-opstring data length, followed by a hidden layer of 32 neurons, and finally an output layer consisting of a few neurons, depending on the number of phases, with softmax activation. In contrast, for regression, the ML model consists of two fully-connected hidden layers, each with $64$ neurons and the final layer has a single neuron with a ReLU activation function.
@@
For phase boundary detection in the 1D SSH model, we have used a CNN architecture. After the input layer, this architecture includes $4$ hidden CNN layers with $16,32,64$ and $128$ filters of sizes $9,5,5$ and $5$, respectively. Each CNN layer is followed by a MaxPool layer with a pool size of $2$ and a stride of $2$, which reduces the length of the 1D data by half. The output of the last CNN layer is then flattened and passed to a fully-connected layer with two neurons, corresponding to the two phases of the SSH model. The ML models used for the phase classification and regression tasks discussed above are schematically represented in Table 1.
@@ IV.1 Two-dimensional $t_{2}-V_{1}$ model
$H=-t_{2}\sum_{\langle\langle i,j\rangle\rangle}(b^{\dagger}_{i}b_{j}+H.c.)+V_{1}\sum_{\langle i,j\rangle}n_{i}n_{j}-\mu\sum_{i}n_{i},$ (8)
@@
Figure 4: Top: ML prediction for the three phases, cS, cSS and MI, as a function of the chemical potential $\mu$. This is for a fixed value of NN repulsion $V_{1}=4$. The vertical dashed lines indicate the ML predicted phase boundaries. Bottom: QMC results for the observables, particle-density $\rho$, superfluid density $\rho_{s}$ and structure factor $S(\pi,\pi)$ with varying chemical potential $\mu$, on a $12\times 12$ square lattice. The left $y$-axis shows the variation of $\rho_{s}$ and $S(\pi,\pi)$, while the right $y$-axis represents the variation of $\rho$.
@@
The study revealed that at half-filling, as the NN repulsion is increased beyond a critical value $V_{c}$, the system undergoes a mixed-order phase transition (displaying the extreme Thouless effect) from a superfluid (SF) phase to a checkerboard solid (cS) phase, where either one of the sub-lattices is completely occupied. For $V>V_{c}$ as the chemical potential $\mu$ is varied, the system goes through two continuous phase transitions. The first transition takes the system from the cS phase to a checkerboard supersolid (cSS) phase, where all the particles occupy a single sub-lattice and hop within it. At the second phase transition, the cSS phase becomes a Mott insulator (MI) at particle-density $\rho=1$, where all the lattice sites are completely filled. At intermediate particle densities between $\rho=1/2$ and $1$, the model displays either a mixed-order or continuous phase transition (separated by a tricritical point) between the SF and cSS phases as a function of $V_{1}$. In Fig. 2 the above discussion is summarized in the phase diagram of the model, obtained from the SSE QMC study on a $24\times 24$ square lattice, depicted in terms of particle-density $\rho$ and NN repulsion $V_{1}$. It should be noted that in the SSE QMC simulations, different values of particle-densities $\rho$ are accessed by tuning the chemical potential $\mu$ of the system. In particular, for the insulating phases, cS and MI, which are present at $\rho=1/2$ and $1$ respectively, one obtains plateaus in the particle-density $\rho$ as a function of $\mu$.
@@
We apply a supervised ML model trained on the QMC spin-opstring data to predict the phase transitions. For the training, we collect $10000$ spin-opstring data in a small window from deep within each of the phases : cS, cSS, SF and MI [^1] In particular for cS and MI phases, deep within the phase precisely means the $\mu$ values around center of the plateau in $\rho$.. The trained ML model is then used to predict unlabeled data along two lines in the phase diagram: firstly, the vertical line at half-filling, where the system undergoes a mixed-order phase transition from SF to cS as a function of $V_{1}$, and secondly, the horizontal line where at a fixed $V_{1}=4.0$ as function of $\mu$, and hence as function of particle-density $\rho$, the system goes through two continuous transitions, from cS to cSS and then cSS to a MI at $\rho=1$. These two lines are indicated by shaded grey bands in Fig. 2. It should be noted that we observe considerable variation in the opstring lengths among different phases. The QMC simulation in MI phase generates the largest opstring lengths, whereas in the SF phase it is the smallest.
@@
Figure 5: Comparison of superfluid density obtained from the ML regression model with the QMC results obtained for a $12\times 12$ square lattice. The error band of the ML data has been obtained from 100 independently trained ML models.
@@
In the top figure of Fig. 3, we show the ML prediction when tested on a dataset obtained for a $12\times 12$ square lattice at half-filling with a varying NN repulsion $V_{1}$. The figure clearly demonstrates a sharp transition between the SF and cS phase displayed by a jump between the extremum probability values of the output neurons. We contrast the ML prediction with the QMC results on the same lattice size, shown in the bottom of Fig. 3, where two observables, superfluid density $\rho_{s}$ and structure factor $S(\pi,\pi)$, are plotted with increasing NN repulsion. For $V_{1}$ below a critical value, the plot indicates the presence of the SF phase, characterized by a near-maximal value of superfluid density and a zero structure factor. In contrast, for $V_{1}$ exceeding this critical value, the system transitions to the cS phase, where $S(\pi,\pi)$ reaches its maximum value, and $\rho_{s}$ drops to zero. This sharp jump in the order parameters to their respective extrema values is a signature of the extreme Thouless effect and this behavior is also reflected in the ML predictions. Notably, the transition point predicted by the ML model matches perfectly with the QMC result, identifying the critical NN repulsion as $V_{1}=2.36\pm 0.04$. It is important to mention here that we found the Batch Normalization layer to be a crucial ingredient in the ML model for making consistent predictions. We provide details regarding this in Appendix A.
@@
Next, in the top panel of Fig. 4, we present the ML predictions for the same system size at fixed NN repulsion $V_{1}=4.0$ as a function of the chemical potential $\mu$. The plot shows the existence of two phase transitions which appear to be continuous, the first one between phases cS and cSS and the second one between cSS and MI. In the bottom figure, we show the same phase transitions demonstrated by the QMC observables $\rho_{s}$ and $S(\pi,\pi)$, along with the particle-density $\rho$ as a function of $\mu$. The figure illustrates the existence of three phases: checkerboard solid at density $\rho=1/2$ where only the structure factor $S(\pi,\pi)$ is non-zero, checkerboard supersolid for intermediate densities where both $S(\pi,\pi)$ and $\rho_{s}$ are non-zero and the Mott insulator at density $\rho=1$ where both the observables, $S(\pi,\pi)$ and $\rho_{s}$, are zero. In the phase diagram in Fig. 2, these two transitions are depicted in terms of the corresponding particle density $\rho$ highlighted by the horizontal shaded band. In this case as well the ML predictions match well with the QMC results.
@@
Finally, we wanted to see how much information can ML extract from our proposed input type in regression tasks. For this we selected the observable superfluid density $\rho_{s}$, which is connected to the winding of particles through the spatial dimensions of the system, thereby capturing non-local properties of the quantum system. The regression task is performed along the same line in the phase diagram as in Fig. 4, with a fixed NN repulsion $V_{1}=4.0$. To train the model we used the same data collected as before, but this time the values of $\rho_{s}$ are used as labels. Since $\rho_{s}$ is an expectation value calculated by averaging over several simulation cells, we have labeled all the spin-opstring data at a given $\mu$ with the same value, i.e., the expectation value $\rho_{s}$ at that $\mu$. Similarly, ML predictions are made for multiple spin-opstring data at a fixed value of $\mu$, and the results are then averaged to obtain the final predicted $\rho_{s}$. In Fig. 5, we compare the prediction of the regression model with the QMC results obtained for a $12\times 12$ square lattice. For the most part the predicted superfluid density values closely agree with the QMC results. However, near the transition points unsurprisingly the machine prediction show deviations as it encounters critical behavior which is absent in the training data. The regression results further support the suitability of our proposed input data for machine learning applications.
@@ V.1 Domain Adaptation
For this purpose we first consider the well-known $t_{1}-V_{1}$ model on the periodic square lattice, governed by the Hamiltonian,
$H=-t_{1}\sum_{\langle i,j\rangle}(b^{\dagger}_{i}b_{j}+H.c.)+V_{1}\sum_{\langle i,j\rangle}n_{i}n_{j}-\mu\sum_{i}n_{i}.$ (12)
Next, we explore the application of transfer learning to the 1D SSH model. We leverage the spin-opstring data to transfer knowledge from the non-interacting SSH model to its interacting version. In the presence of NN repulsion, the Hamiltonian of the interacting SSH model reads as,
@@
where $H_{\mathrm{NI}}$ corresponds to the non-interacting SSH Hamiltonian given in Eq. 11 and $\langle i,j\rangle$ denotes NN pairs of sites. Even with the inclusion of NN repulsion, the topological phase of the SSH model persists up to a certain interaction strength, with the transition to the trivial phase occurring at $v=w$. However, for larger values of $V_{1}$, a charge-density-wave state emerges between the topological and trivial phases[35, 36, 37].
@@
In our transfer learning setup, we trained this CNN model exclusively on data from the non-interacting SSH regime and tested its performance on the interacting SSH model. Fig. 10 shows the transition points between the topological and trivial phases predicted by the ML model, as a function of increasing repulsion $V_{1}$. It demonstrates that the model is capable of transferring the knowledge learned in the non-interacting limit to accurately predict the transition points for the interacting system up to $V_{1}\approx 2$. The model delivers inconsistent predictions for larger repulsions as the distribution of states differs significantly from the non-interacting system, in particular, due to the emergence of the CDW phase for large $V_{1}$.
Prior work has demonstrated transfer learning in the context of SSH model by training a ML model on data from a clean system and applying it to a disordered one [38]. The eigenstates of the SSH Hamiltonian, under open and periodic boundary conditions, were used as the ML input. It was observed that standard CNN models were unable to generalize effectively in the presence of disorder. With the application of a domain adversarial neural network, this challenge was overcome for the PBC case, while for data with OBC, the incorporation of unsupervised techniques was found to be necessary. Although our focus is on a different domain adaptation task, it is worth mentioning that simple CNN models, trained with spin-opstring data obtained using PBC, were successful in transfer learning.

### v3

@@ Abstract
We propose the use of the “spin-opstring”, derived from Stochastic Series Expansion Quantum Monte Carlo (QMC) simulations as machine learning (ML) input data. It offers a compact, memory-efficient representation of QMC simulation cells, combining the initial state with an operator string that encodes the state’s evolution through imaginary time. Using supervised ML, we demonstrate the input’s effectiveness in capturing both conventional and topological phase transitions, and in a regression task to predict non-local observables. We also demonstrate the capability of spin-opstring data in transfer learning by training models on one quantum system and successfully predicting on another, as well as showing that models trained on smaller system sizes generalize well to larger ones. Importantly, we illustrate a clear advantage of spin-opstring over conventional spin configurations in the accurate prediction of a quantum phase transition. Finally, we show how the inherent structure of spin-opstring provides an elegant framework for the interpretability of ML predictions. Using two state-of-the-art interpretability techniques, Layer-wise Relevance Propagation and SHapley Additive exPlanations, we show that the ML models learn and rely on physically meaningful features from the input data. Together, these findings establish the spin-opstring as a broadly-applicable and interpretable input format for ML in quantum many-body physics.
@@ I Introduction
We further explored the application of spin-opstring data for transfer learning tasks. First, we were able to predict phase transitions in the interacting SSH model using a ML model trained in the non-interacting case. Next, we applied our ML model trained on data from the $t_{2}-V_{1}$ model to successfully predict a phase transition in the $t_{1}-V_{1}$ model. Additionally, we showed that it is possible to generalize a ML model trained with spin-opstring data from a smaller system to larger systems.
With the spin-opstring representation proving effective across various ML tasks, we investigate its potential advantages over other data formats, such as spin configurations. To explore this, we study a two-dimensional hard-core boson model that hosts a weak topological insulator phase. We show that spin-opstring outperforms spin-state due to the additional quantum information it encodes. Additionally, we demonstrate that it also surpasses the operator string alone, highlighting the necessity of combining both components into a single representation.
Having established the usefulness of this input datatype, we turn our focus to the interpretation of the decision-making process in the trained ML models. It is important to mention here that the majority of studies dealing with interpretability put their emphasis on spin states as the input data [24, 25, 26]. The suitability of alternative input representations in such analyses remain mostly unexplored [27, Cybinski_2025]. To the best of our knowledge, our study marks the first instance of interpreting ML models trained on simulation cell data. The structure of the spin-opstring, composed of a spin-state and a sequence of operators, provides an elegant framework for interpreting ML models. To this end, we utilize two well-known interpretability methods, Layer-wise Relevance Propagation (LRP) [29] and SHapley Additive exPlanations (SHAP) [30] to understand which features in the data are most important to the model’s predictions. We developed a strategy to extract meaningful insights, based on different physical operators, from the large-dimensional feature scores obtained directly from LRP and SHAP. In the classification, regression and transfer learning tasks, we find that ML models trained with spin-opstring data are able to identify key physical features from each phase, showcasing their ability to learn and utilize essential quantum characteristics.
@@ II Proposal
Figure 1: Schematic diagram of the simulation cell of SSE QMC algorithm for a spin-$1/2$ system with six sites. Here $i$ signifies the lattice site number and $\sigma(i)$ represents the initial/final spin state. At each level $p$ of the simulation cell, the spin state is depicted by filled (+1)/open (-1) circles. The diagonal or off-diagonal operator at each level acting on a bond is represented by open or filled bars respectively. The information in the operator string is compactly codified in $S(p)$ ($a(p)$ and $b(p)$ are defined in the text). The “spin-opstring” for the above simulation cell is constructed by combining $\sigma(i)$ and $S(p)$ together.
@@
$|\alpha(p)\rangle=\prod\limits_{i=1}^{p}H_{a(i),b(i)}|\alpha\rangle,\quad|\alpha(0)\rangle=|\alpha\rangle,$ (7)
@@ III Method
For the phase classification in the 2D $t_{2}-V_{1}$ model, we have used a FC ML model with three layers: the first input layer consisting of a number of neurons equal to the spin-opstring data length, followed by a hidden layer of 32 neurons, and finally an output layer consisting of a few neurons, depending on the number of phases, with softmax activation. In contrast, for regression, the ML model consists of two fully-connected hidden layers, each with $64$ neurons and the final layer has a single neuron with a ReLU activation function.
@@
For phase boundary detection in the 1D SSH model, we have used a CNN architecture. After the input layer, this architecture includes $4$ hidden CNN layers with $16,32,64$ and $128$ filters of sizes $9,5,5$ and $5$, respectively. Each CNN layer is followed by a MaxPool layer with a pool size of $2$ and a stride of $2$, which reduces the length of the 1D data by half. The output of the last CNN layer is then flattened and passed to a FC layer with two neurons, corresponding to the two phases of the SSH model. The ML models used for the phase classification and regression tasks discussed above are schematically represented in Table 1.
@@ IV.1 Two-dimensional $t_{2}-V_{1}$ model
$H=-t_{2}\sum_{\langle\langle i,j\rangle\rangle}(b^{\dagger}_{i}b_{j}+\textrm{h.c.})+V_{1}\sum_{\langle i,j\rangle}n_{i}n_{j}-\mu\sum_{i}n_{i},$ (8)
@@
Figure 4: Top: ML prediction for the three phases, cS, cSS and MI, as a function of the chemical potential $\mu$ for a fixed NN repulsion $V_{1}=4$. The vertical dashed lines indicate the ML predicted phase boundaries. Bottom: QMC results for the observables, particle-density $\rho$, superfluid density $\rho_{s}$ and structure factor $S(\pi,\pi)$ with varying chemical potential $\mu$, on a $12\times 12$ square lattice. The left $y$-axis shows the variation of $\rho_{s}$ and $S(\pi,\pi)$, while the right $y$-axis represents the variation of $\rho$.
@@
The study revealed that at half-filling, as the NN repulsion is increased beyond a critical value $V_{1}^{c}$, the system undergoes a mixed-order phase transition (displaying the extreme Thouless effect) from a superfluid (SF) phase to a checkerboard solid (cS) phase, where either one of the sub-lattices is completely occupied. For $V_{1}>V_{1}^{c}$ as the chemical potential $\mu$ is varied, the system goes through two continuous phase transitions. The first transition takes the system from the cS phase to a checkerboard supersolid (cSS) phase, where all the particles occupy a single sub-lattice and hop within it. At the second phase transition, the cSS phase becomes a Mott insulator (MI) at particle-density $\rho=1$, where all the lattice sites are completely filled. At intermediate particle densities between $\rho=1/2$ and $1$, the model displays either a mixed-order or continuous phase transition (separated by a tricritical point) between the SF and cSS phases as a function of $V_{1}$. In Fig. 2 the above discussion is summarized in the phase diagram of the model, obtained from the SSE QMC study on a $24\times 24$ square lattice, depicted in terms of particle-density $\rho$ and NN repulsion $V_{1}$. It should be noted that in the SSE QMC simulations, different values of particle-densities $\rho$ are accessed by tuning the chemical potential $\mu$ of the system. In particular, for the insulating phases, cS and MI, which are present at $\rho=1/2$ and $1$ respectively, one obtains plateaus in the particle-density $\rho$ as a function of $\mu$.
@@
We apply a supervised ML model trained on the QMC spin-opstring data to predict the phase transitions. For the training, we collect $10000$ spin-opstring data in a small window from deep within each of the phases : cS, cSS, SF and MI [^1] In particular for cS and MI phases, deep within the phase precisely means the $\mu$ values around center of the plateau in $\rho$.. The trained ML model is then used to predict unlabeled data along two lines in the phase diagram: firstly, the vertical line at half-filling, as a function of $V_{1}$, and secondly, the horizontal line at a fixed $V_{1}=4.0$ as function of $\mu$, and hence as function of particle-density $\rho$. These two lines are indicated by shaded grey bands in Fig. 2. Notably, we observe considerable variation in opstring lengths across phases, with the longest lengths occurring in the MI phase and the shortest in the SF phase.
@@
Figure 5: Comparison of superfluid density obtained from the ML regression model with QMC results on a $12\times 12$ square lattice. The error band of the ML data has been obtained from 100 independently trained ML models.
@@
In Fig. 3 top, we show the ML prediction on a $12\times 12$ square lattice at half-filling with a varying NN repulsion $V_{1}$. A sharp transition between the SF and cS phase is displayed by a jump between the extremum probability values of the output neurons. We contrast the ML prediction with the QMC results on the same lattice size (Fig. 3 bottom), where superfluid density $\rho_{s}$ and structure factor $S(\pi,\pi)$, are plotted with increasing $V_{1}$. For $V_{1}$ below a critical value, the presence of the SF phase is characterized by a near-maximal value of $\rho_{s}$ and a zero $S(\pi,\pi)$. With increasing $V_{1}$, the system transitions to the cS phase, where $S(\pi,\pi)$ reaches its maximum value, and $\rho_{s}$ drops to zero. This sharp jump in the order parameters to their respective extrema values is a signature of the extreme Thouless effect and this behavior is also reflected in the ML predictions. Notably, the transition point predicted by the ML model matches perfectly with the QMC result, identifying the critical NN repulsion as $V_{1}=2.36\pm 0.04$. It is important to mention here that we found the Batch Normalization layer to be a crucial ingredient in the ML model for making consistent predictions. We provide details regarding this in Appendix A.
@@
Next, in Fig. 4 top, we present the ML predictions for the same system size at fixed $V_{1}=4.0$ as a function of $\mu$, revealing two phase transitions, from cS to cSS, and from cSS to MI, both of which appear to be continuous. In the bottom figure, we show the same phase transitions demonstrated by the QMC observables $\rho_{s}$ and $S(\pi,\pi)$, along with the particle-density $\rho$ as a function of $\mu$. The figure illustrates the existence of three phases: checkerboard solid at density $\rho=1/2$ where only the structure factor $S(\pi,\pi)$ is non-zero, checkerboard supersolid for intermediate densities where both $S(\pi,\pi)$ and $\rho_{s}$ are non-zero and the Mott insulator at density $\rho=1$ where both the observables, $S(\pi,\pi)$ and $\rho_{s}$, are zero. In the phase diagram in Fig. 2, these two transitions are depicted in terms of the corresponding particle density $\rho$ highlighted by the horizontal shaded band. In this case as well the ML predictions match well with the QMC results.
@@
Next, to evaluate the performance of the spin-opstring data type in regression tasks, we selected superfluid density $\rho_{s}$, a non-local observable connected to the winding of particles through the spatial dimensions of the system. The regression task is performed along the same line as in Fig. 4, with $V_{1}=4.0$, using the same data, now labeled with the values of $\rho_{s}$. Since $\rho_{s}$ is an expectation value calculated by averaging over several simulation cells, all samples at a given $\mu$ share the same $\rho_{s}$ value as label. Similarly, ML predictions are averaged over multiple samples for each $\mu$ to obtain the final prediction for $\rho_{s}$. In Fig. 5, we compare the prediction of the regression model with the QMC results obtained for a $12\times 12$ square lattice. For the most part the predicted superfluid density values closely agree with the QMC results. However, near the transition points unsurprisingly the machine prediction show deviations as it encounters critical behavior which is absent in the training data. The regression results further support the suitability of our proposed input data for machine learning applications.
@@ V.1 Domain Adaptation
For this purpose we consider the 1D SSH model. We leverage the spin-opstring data to transfer knowledge from the non-interacting SSH model to its interacting version. In the presence of NN repulsion, the Hamiltonian of the interacting SSH model reads as,
@@
where $H_{\mathrm{NI}}$ corresponds to the non-interacting SSH Hamiltonian given in Eq. 11 and $\langle i,j\rangle$ denotes NN pairs of sites. Even with the inclusion of NN repulsion, the topological phase of the SSH model persists up to a certain interaction strength, with the transition to the trivial phase occurring at $v=w$. However, for larger values of $V_{1}$, a charge-density-wave (CDW) state emerges between the topological and trivial phases[35, 36, 37].
@@
In our transfer learning setup, we trained a CNN model exclusively on data from the non-interacting SSH regime and tested its performance on the interacting SSH model. Fig. 9 shows the transition points between the topological and trivial phases predicted by the ML model, as a function of increasing repulsion $V_{1}$. It demonstrates that the model is capable of transferring the knowledge learned in the non-interacting limit to accurately predict the transition points for the interacting system up to $V_{1}\approx 2$. The model delivers inconsistent predictions for larger repulsions as the distribution of states differs significantly from the non-interacting system, in particular, due to the emergence of the CDW phase for large $V_{1}$.
An additional example of domain adaptation, where a ML model trained on data from the $t_{2}-V_{1}$ model successfully predicts a phase transition in the $t_{1}-V_{1}$ model, is provided in Appendix B.
In the context of domain adaptation in 1D SSH model, Ref. [38] employed a supervised ML algorithm to identify topological phases in interacting systems by training a neural network on their non-interacting versions, with the curvature function at the high symmetry points as input. Another study [39] also demonstrated transfer learning in the SSH model, from a clean to disordered system, using eigenstates as input with domain adversarial neural networks.